74=(69+75+60+x)/4
Show steps to solve

Answer by Jstrasner(112)

Hey,
So for this problem we need to isolate the x:
First we need to get rid of the 4 in the denominator on the right side by multiplying 4 on both sides so it will cancel out on the right side:
4 x 74 = (69+75+60+x)/4 x 4 => 296 = 69 + 75 + 60 + x
Now we need to add up all the numbers on the right side and subtract that number from both sides so that the x is left standing by itself:
296 = 240 + x => 296 - 240 = 240 + x - 240 => 56 = x Therefore:
x = 56
I hope this helps!